OpenStudy (anonymous):

if 9-5x=12-(6x+7) evaluate x(second power)-3x-2

OpenStudy (sandra):

so one way to do this is to solve for x 1. 9 - 5x = 12 - (6x + 7) 2. 9 - 5x = 12 -6x - 7 3. 9 - 5x = 5 - 6x 4. 4 - 5x = -6x 5. 4 = -x 6. x = -4

OpenStudy (sandra):

so now that you have the value of x, you can stick it into your second equation

OpenStudy (sandra):

does that make sense?

OpenStudy (anonymous):

oh ok so i had to solve that one part to evaluate the other part

OpenStudy (sandra):

right exactly

OpenStudy (sandra):

this is probably setting you up to learn how to do substitution problems, in which the last step is usually inserting the solution to one equation into the next

OpenStudy (sandra):

(subsituting)

OpenStudy (anonymous):

k thanks sandra!